(2X-3)/(X+1) less than or equal to 1
Answer by stanbon(75887)

(2X-3)/(X+1) <= 1
----------------------
Draw a number line and plot x = -1 and x = 3/2
------------------------------------
For x < -1, x+1 < 0
Then (2x-3) >= x+1
x >= 4
But x cannot be <-1 and also >= 4
----
For x > -1
2x-3 <= x+1
x <= 4
---
Final Solution: -1 < x <= 4
======================================
Cheers,
Stan H.